# Student Member, IEEE

### v1

@@ 7 Model Predictive Controller Objective
$\quad\;\;\!\!\sum_{t=k}^{k+K+1}\!\!\gamma^{t-k}\!\!\sum_{m=1}^{M}\!\sum_{n=1}^{N}\!\mathbf{1}\!\left(\lVert\bm{\chi}_{t}^{T_{n}}\!+\!\bm{\chi}_{t}^{R_{m}}(\bm{u}_{t})\rVert\!\leq\!R_{R2T}\!\right)$
@@ 7.1 Model Predictive Path Integral
where $P(\cdot)$ is the PDF of an control input sequence in an uncontrolled system ($\bm{u}_{k:k+K+1}=0)$, $b$ is the temperature of the distribution, and $\eta$ is the evidence. However, the optimal control PDF likely does not possess an analytical form. Therefore, by minimizing the Kullback-Leibler Divergence (KL-Diverence) betwewen $Q^{*}$ and a variational Gaussian PDF $Q$ parameterized by $u,\Sigma$ [24] with respect to $u$ gives the optimal control at $k$ as [24]
@@
Due to the high dimensionality of the targets’ control input over a horizon and the limited number of samples in MPPI, the sample covariance matrix for control inputs can overfit, leading to inaccurate variances [50]. This often results in poor control inputs from the proposal distribution of MPPI. To address this, we adopt the Ledoit-Wolf (LW) for AIS, which approximates Oracle Approximating Shrinkage (OAS) shrinkage and helps mitigate overfitting, following the method described in [25].
@@ 8 EXPERIMENTS
We examine three scenarios: 1. an “underdetermined” system, with fewer radars than targets, 2. an “overdetermined” system, with more radars than targets, and 3. an “full” system, with equal amount of radar and targets. We highlight not only the poorer performance of a misspecified measurement model (DDR versus constant covariance range (CCR) measurement model) but also the pathologies in radar movement caused by the controller using a FIM-based objective specified from the CCR measurement model. The proposed approach outperforms methods employing stationary radars and simplified range measurement models in target localization, achieving a reduction in mean RMSE and the upper tail of the 90% Highest Density Interval (HDI) by 38-74% and 33-79% respectively, across 500 MC trials at all time steps.
@@ 8.1 Simulation
The hyperparameters for each experiment for the MPPI control, the AIS, the radar configurations , and the MPC objective is found in Tables 3, 2 and 1. We assume that the radars move on the ground plane (z=0), while the drones move in free space by the constant velocity kinematic model (Equation 5) with acceleration noise constant $\sigma_{W}=\sqrt{10}$. For each experiment , we conduct 500 MC trials, where every MC trial is 600 time steps (1 minute in real time). The initial heading velocity and acceleration of the radars are both set to 0, and their initial positions in the XY-plane are uniformly sampled within a square with edge sizes 800 [m], centered at the origin (0 [m], 0 [m]). Although the targets may start in the same position for each MC trial, the random acceleration noise in the transition model (Equation 8) leads to significant variations in the trajectories of the targets across trials. The observed range measurements at each time step are simulated using the more realistic range measurement model presented in Equation 4.
$\sigma_{a}^{2}=\frac{M\tilde{P_{r}}}{10^{SNR/10}}$ (38)
8.2 Improved Target Localization
In over 500 MC trials, we empirically show that our approach exhibits significantly lower RMSE for all time steps (see Figure 3 (a,b,d,e,g,h)) compared to both the stationary radar approach and the mobile radar approach with constant range measurement. Additionally, the 90% HDI interval in localization performance for the CCR measurement model is significantly larger compared to the DDR measurement model presented in [16] (Figure 3 (c,f,i)). The PFIM-based objective for both range measurement models has slightly higher targets’ state estimate RMSE compared to the SFIM-based objective. However, Bayesian methods are notorious for overfitting, attributed to ensembling with model mispecification and the assumption of a single true model generating the data [51, 52]. We further examine the performance of the proposed approach inn the three simulation scenarios in the next subsections.
8.2.1 3 Mobile Radars and 4 Targets
We initialize the targets such that all target directional velocities point towards the upper right (northeast) in the world coordinate frame, and begin around the same position (see Figure 2. Further details about the initial target state is found in Subsection .2). Therefore, with fewer radars than targets, we anticipate a circular formation of radars near the targets when continuously optimizing radar placement, aiming to maintain nearest feasible proximity to the targets.
Figure 4: RMSE for single simulation realization in Figure 2
8.2.2 6 Mobile Radars and 3 Targets
We initialize each target’s velocity direction to be different, with target 1 heading towards the upper right, target 2 towards the lower right, and target 3 towards the lower left. The initial positions of the targets are depicted in Figure 2. Further details are available in Subsection .2. Since the targets’ are spreading out over an open space, we hypothesize that the radars will send atleast 1 radar to follow each target, while keeping a reasonable distance to all other targets.
Figure 5: RMSE for single simulation realization in Figure 2
8.2.3 4 Mobile Radars and 4 Targets
We initialize two targets to move towards the upper right and two targets to move towards the lower left. The initial positions of the targets are shown in Figure 3. Further details can be found in Subsection .2. We hypothesize that the radars will pair off, with two radars following each group of two targets.
Figure 6: RMSE for single simulation realization in Figure 3
8.2.4 Stationary Radars
9 CONCLUSION
We implemented two improvements to optimize radar placement for multi-target localization: a more realistic radar range measurement model and an efficient stochastic optimal controller (Model Predictive Path Integral (MPPI)) for handling non-smooth, non-convex objectives in real time. Our approach significantly improves the Cubature Kalman Filter (CKF) target state estimate Root Mean Squared Error (RMSE) compared to stationary radars. Additionally, we show the pitfalls of the constant covariance range (CCR) measurement model, which may increase target localization error by moving radars in the opposite direction of target movement to optimize angles rather than the radar-target distance. Overall, the proposed strategy outperforms stationary radars and simplified range measurement models in target localization, achieving a 38-74% reduction in mean RMSE and an 33-79% reduction in the upper tail of the 90% Highest Density Interval (HDI) over 500 Monte Carlo (MC) trials across all time steps. For future work, we plan to integrate occupancy grid maps into the collision cost function to account for static and dynamic obstacles, such as trees and buildings, and further improve the radar range measurement model with traditionally Multi-Target Tracking (MTT) characteristics.
Appendix Table 1: The MPPI configuration settings for each experiment
@@ 9 CONCLUSION
Appendix Table 2: The MPPI settings for each experiment
@@
Appendix Table 3: The MPC configuration settings for each experiment
.2.1 3 Radars and 4 Targets Initial State
The initial targets’ state is
@@
.2.2 6 Radars and 3 Targets Initial State
The initial targets’ state is
@@
.2.3 4 Radars and 4 Targets Initial State
The initial targets’ state is
@@
.3 Single Simulation Realization
The extended MC simulation over 1000 time steps for 6 radars and 3 targets, and 4 radars and 4 targets are shown in Figure 2 and Figure 3, respectively. GIFs for the undetermined, overdetermined, and full system are included in supplemental materials.
(a) Proposed approach: Each radar moves towards target
(b) [9, 13, 22] range model: Radars move to encircle targets
(c) Radars move toward targets as closely as feasible
(d) Radars encircle targets
(e) Two radars follow each target individually
(f) Radars encircle the targets, but the upper most two radars move far away from the upper most target.
(g) Two radars track each target individually by moving side-by-side parallel to the target’s trajectory.
(h) Radars maintain encirclement formation around target, but the uppermost radar is far from the uppermost target
Appendix Figure 2: 6 radars and 3 targets. The left column shows snapshots of target localization using a range measurement model with covariance dependent on radar-target distance [16], while the right column depicts tracking with a range measurement model employing constant diagonal covariance [13]. In the first three rows, each subfigure’s left panel shows a 3D projection onto 2D of the radar placement at a specific time. The colored lines emanating from the radars depict the weighted MPPI-planned trajectories, with blue indicating better paths and red indicating worse ones. Dashed red and green circles mark the radar’s (10 [m]) and the target’s (125 [m]) collision avoidance boundaries, respectively, while blue dashed circles represent the targets’ collision avoidance boundary based on the CKF position estimate. The radar’s trajectory is the last 25 time steps. The right panel displays the log determinant of the FIM (for the ddr measurement model) concerning the 2D target position for the specific time step, meshed over the free space area, and evaluated at each mesh point.
(a) Proposed approach: Pairs of radars move towards pairs of targets
(b) [9, 13, 22] range model: the upperleft radar moves away from target
(c) Radars move toward targets as closely as feasible, with pairs of radars positioned side by side with pairs of targets.
(d) Radars encircle the targets, with the upper left radar positioned farther away from the targets.
(e) The pairs of targets are kept in the maximum information gain locations of the radar pairs.
(f) The bottom left pair of targets only has one radar following, leading to worse target localization.
(g) Pairs of radars track each pair of targets by moving side-by-side parallel to the pair of targets’ trajectory.
(h) Radars maintain encirclement formation around target, but the bottom left radar is far from the bottom left target
Appendix Figure 3: 4 radars and 4 targets. The left column shows snapshots of target localization using a range measurement model with covariance dependent on radar-target distance [16], while the right column depicts tracking with a range measurement model employing constant diagonal covariance [13]. In the first three rows, each subfigure’s left panel shows a 3D projection onto 2D of the radar placement at a specific time. The colored lines emanating from the radars depict the weighted MPPI-planned trajectories, with blue indicating better paths and red indicating worse ones. Dashed red and green circles mark the radar’s (10 [m]) and the target’s (125 [m]) collision avoidance boundaries, respectively, while blue dashed circles represent the targets’ collision avoidance boundary based on the CKF position estimate. The radar’s trajectory is the last 25 time steps. The right panel displays the log determinant of the FIM (for the ddr measurement model) concerning the 2D target position for the specific time step, meshed over the free space area, and evaluated at each mesh point.

### v3

@@ 7 Model Predictive Controller Objective
$\quad\;\;\!\!\sum_{t=k}^{k+K+1}\!\!\gamma^{t-k}\!\!\sum_{m=1}^{M}\!\sum_{n=1}^{N}\!\mathbf{1}\!\left(\lVert\bm{\chi}_{t}^{T_{n}}\!-\!\bm{\chi}_{t}^{R_{m}}(\bm{u}_{t})\rVert\!\leq\!R_{R2T}\!\right)$
@@ 7.1 Model Predictive Path Integral
where $P(\cdot)$ is the PDF of an control input sequence in an uncontrolled system ($\bm{u}_{k:k+K+1}=0)$, $b$ is the temperature of the distribution, and $\eta$ is the evidence. However, the optimal control PDF likely does not possess an analytical form. Therefore, by minimizing the Kullback-Leibler Divergence (KL-Divergence) between $Q^{*}$ and a variational Gaussian PDF $Q$ parameterized by $u,\Sigma$ [24] with respect to $u$ gives the optimal control at $k$ as [24]
@@
Due to the high dimensionality of the targets’ control input over a horizon and the limited number of samples in MPPI, the sample covariance matrix for control inputs can overfit, leading to inaccurate variances [51]. This often results in poor control inputs from the proposal distribution of MPPI. To address this, we adopt the Ledoit-Wolf (LW) for AIS, which approximates Oracle Approximating Shrinkage (OAS) shrinkage and helps mitigate overfitting, following the method described in [25]  [^1] The authors in [25] do not explicitly discuss Oracle Approximating Shrinkage (OAS) or Ledoit-Wolfe, but their publicly available code demonstrates methods for estimating the sample covariance, including Ledoit & Wolf, Schaffer & Strimmer, Rao-Blackwell, and OAS. A reference in the publicly available code is https://github.com/sisl/MPOPIS/wiki/MPOPIS-Details..
@@
(a) 3 Radar - 4 Target: RMSE with 90% HDI
(b) 3 Radar - 4 Target: RMSE without stationary radar or HDI
(c) 3 Radar - 4 Target: RMSE ECDF
(d) 6 Radar - 3 Target: RMSE with 90% HDI
(e) 6 Radar - 3 Target: RMSE without stationary radar or HDI
(f) 6 Radar - 3 Target: RMSE ECDF
(g) 4 Radar - 4 Target: RMSE with 90% HDI
(h) 4 Radar - 4 Target: RMSE without stationary radar or HDI
(i) 4 Radar - 4 Target: RMSE ECDF
@@ 8 EXPERIMENTS
We examine three scenarios: 1. an “underdetermined” system, with fewer radars than targets, 2. an “overdetermined” system, with more radars than targets, and 3. an “full” system, with equal amount of radar and targets. We highlight not only the poorer performance of a misspecified measurement model (DDR versus constant covariance range (CCR) measurement model) but also the pathologies in radar movement caused by the controller using a FIM-based objective specified from the CCR measurement model. The proposed approach outperforms methods employing stationary radars and simplified range measurement models in target localization, achieving a reduction in mean RMSE and the upper tail of the 90% HDI by 38-74% and 33-79% respectively, across 500 Monte Carlo (MC) trials at all time steps.
@@ 8.1 Simulation
The hyperparameters for each experiment for the MPPI control, the AIS, the radar configurations , and the MPC objective is found in Tables 3, 2 and 1.
Radar Parameter Value
Carrier Frequency $1\times 10^{8}$
Radar Loss 1
Transmitter Gain 200
Receiver Gain 200
Radius (for SNR) 500 [m]
Radar Cross Section 1 [m2]
Power Transmit 1000 [W]
SNR -20 [dB]
Table 1: The Radar Equation configuration settings for each experiment
@@
Table 2: The MPPI settings for each experiment
@@
Table 3: The MPC configuration settings for each experiment
We assume that the radars move on the ground plane (z=0), while the drones move in free space by the constant velocity kinematic model (Equation 5) with acceleration noise constant $\sigma_{W}=\sqrt{10}$. For each experiment , we conduct 500 MC trials, where every MC trial is 600 time steps (1 minute in real time). The initial heading velocity and acceleration of the radars are both set to 0, and their initial positions in the XY-plane are uniformly sampled within a square with edge sizes 800 [m], centered at the origin (0 [m], 0 [m]). Although the targets may start in the same position for each MC trial, the random acceleration noise in the transition model (Equation 9) leads to significant variations in the trajectories of the targets across trials. The observed range measurements at each time step are simulated using the more realistic range measurement model presented in Equation 4.
$\sigma_{a}^{2}=\frac{M\tilde{P_{r}}}{10^{SNR/10}}$ (39)
8.2 Improved Target Localization
In over 500 MC trials, we empirically show that our approach exhibits significantly lower RMSE for all time steps (see Figure 4 (a,b,d,e,g,h)) compared to both the stationary radar approach and the mobile radar approach with constant range measurement. Additionally, the 90% HDI interval in localization performance for the CCR measurement model is significantly larger compared to the DDR measurement model presented in [16] (Figure 4 (c,f,i)). The PFIM-based objective for both range measurement models has slightly higher targets’ state estimate RMSE compared to the SFIM-based objective.
However, Bayesian methods are notorious for overfitting, attributed to ensembling with model mispecification and the assumption of a single true model generating the data [52, 53]. We further examine the performance of the proposed approach inn the three simulation scenarios in the next subsections.
8.2.1 3 Mobile Radars and 4 Targets
We initialize the targets such that all target directional velocities point towards the upper right (northeast) in the world coordinate frame, and begin around the same position (see Figure 3):
@@ 8.2 Improved Target Localization
Therefore, with fewer radars than targets, we anticipate a circular formation of radars near the targets when continuously optimizing radar placement, aiming to maintain nearest feasible proximity to the targets.
Figure 5: RMSE for single simulation realization in Figure 3
8.2.2 6 Mobile Radars and 3 Targets
We initialize each target’s velocity direction to be different, with target 1 heading towards the upper right, target 2 towards the lower right, and target 3 towards the lower left:
@@
The initial positions of the targets are depicted in Figure 1. Since the targets’ are spreading out over an open space, we hypothesize that the radars will send atleast 1 radar to follow each target, while keeping a reasonable distance to all other targets.
Figure 6: RMSE for single simulation realization in Figure 1
8.2.3 4 Mobile Radars and 4 Targets
We initialize two targets to move towards the upper right and two targets to move towards the lower left. The initial positions of the targets are:
@@
We hypothesize that the radars will pair off, with two radars following each group of two targets.
Figure 7: RMSE for single simulation realization in Figure 2
8.2.4 Stationary Radars
9 CONCLUSION
We implemented two improvements to optimize radar placement for multi-target localization: a more realistic radar range measurement model and an efficient stochastic optimal controller (Model Predictive Path Integral (MPPI)) for handling non-smooth, non-convex objectives in real time. Our approach significantly improves the Cubature Kalman Filter (CKF) target state estimate Root Mean Squared Error (RMSE) compared to stationary radars. Additionally, we show the pitfalls of the constant covariance range (CCR) measurement model, which may increase target localization error by moving radars in the opposite direction of target movement to optimize angles rather than the radar-target distance. Overall, the proposed strategy outperforms stationary radars and simplified range measurement models in target localization, achieving a 38-74% reduction in mean RMSE and an 33-79% reduction in the upper tail of the 90% Highest Density Interval (HDI) over 500 Monte Carlo (MC) trials across all time steps. For future work, we plan to integrate occupancy grid maps into the collision cost function to account for static and dynamic obstacles, such as trees and buildings, and further improve the radar range measurement model with traditional Multi-Target Tracking (MTT) characteristics.